x=y-4
2x+8=2y
solve using elimination, substitution or graphing
could this be no solutions?

Answer by checkley71(8403)

X=Y-4 THEN 2(Y-4)+8=2Y OR 2Y-8+8=2Y OR 2Y-2Y=8-8 OR 0=0 NO SOLUTION